OpenStudy (anonymous):

Adam has 4 dimes, 4 quarters and 4 nickels in his pocket. He pulls out 2 coins making sure to put the first one back in his pocket before picking the second. What are the odds that he pulls out 2 nickels?

OpenStudy (cwrw238):

don't just give the answer princess explain how you got it

OpenStudy (cwrw238):

her answer makes no sense anyway he has a total of 12 coins in his pocket of which 4 are nickels so what are the odds he pols out a nickel?

OpenStudy (anonymous):

You have to multiply P(getting one nickel)*P(getting one nickel) where P(getting one nickel) = (4 nickels)/(4 nickels+4 dimes +4 quarters)

OpenStudy (anonymous):

4/12 ?

OpenStudy (cwrw238):

right so P( 1 nickel) = 4/12 = 1/3

OpenStudy (anonymous):

thank you bro

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

though i'm female! P(another nickel) is also 1/3

OpenStudy (anonymous):

thank you madam :p

OpenStudy (cwrw238):

the 2 picks are independent events and are therefore multiplied together so the final answer is 1/3 * 1/3

OpenStudy (anonymous):

so its 1/9 ?